OpenStudy (anonymous):

Given f(x) = 5 x^2 - 15 x + ( -2 ) , at which point will the curve have slope (a) 0 (b) 2 (c) - 2 All the answers MUST be points. (a) The point where the slope is 0 is . (b) The point where the slope is 2 is . (c) The point where the slope is - 2 is .

OpenStudy (anonymous):

i think this is right

OpenStudy (anonymous):

Juanita, I think they meant the slope was those values, not x..

OpenStudy (anonymous):

dy/dx = 10x-15 gradient= 0, dy/dx= 0 10x-15=0 x=1.5 when x=1.5, f(1.5)= 5(1.5)^2 -15(1.5)-2= -13.25 (1.5, -13.25) dy/dx = 2, 10x-15=2 x=1.7 f(1.7)= -13.05 (1.7,-13.05) dy/dx=-2 10x-15=-2 x=1.3 f(1.3)=-13.05 (1.3,-13.05)

OpenStudy (anonymous):

acekly was correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:D